THEORETICAL STUDIES OF THE FUNDAMENTAL AND OVERTONE SPECTRUM OF OXALIC ACID

\underline{DA MATTHEWS}, JF STANTON, Department of Chemistry and Biochemistry, University of Texas at Austin, Austin, TX 78712.

The vibrational frequencies and intensities for all states up to two quanta are calculated for oxalic acid (C_{2}H_{2}O_{4}) at the CCSD(T) level. The effect of different internal hydrogen bonding configurations on the OH stretching level structures for the five conformers studied is discussed.